OpenStudy (anonymous):

For the function y=x^3-6x^2+9x+1 , find the stationary points and classify their nature. "help"

OpenStudy (amistre64):

stationary points?

OpenStudy (anonymous):

did you mean maxima and minima?

OpenStudy (amistre64):

i gots no idea what a stationary point it, but I would assume that their nature is to be stationary ..

OpenStudy (ash2326):

Stationary points are points where derivative is zero or slope is zero find dy/dx dy/dx=3x^2-12x+9 equate it to 0 3x^2-12x+9=0 x^2-4x+3=0 (x-1)(x-3)=0 x= 1 and x=3 are the two stationary points

OpenStudy (ash2326):

to classify them , let's find d^2y/dx^2 it comes out to be 6x-12 if d^2y/dx^2>0 it is a minima, if it's less than 0 then it is a point of maxima x=1 d^2y/dx^2=-6 so it's a point of maxima x=3 d^2y/dx^2=6 so it's a point of minima

OpenStudy (mattfeury):

Jacob, to respond to people, enter your response here in the reply field. Clicking 'Report Abuse' is for reporting inappropriate content or spam.

OpenStudy (anonymous):

Sorry !

OpenStudy (mattfeury):

No problem :).